Shift report 20010801 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Wed Aug 01 2001 - 15:30:39 EDT

  • Next message: Apache: "Shift report 20010801 16:00-24:00"

    ------------------------------------------------------------
    Date: Wed Aug  1 15:30:38 2001, Shift: 08:00-16:00
    Supervisor: EJK, with: TL,AH
    Log book#6 pages 141 - 142
    
    Shift Summary:
    ------------------------------------------------------------
    1.Had no beam during this shift
    
    2.#3954 Pedestal run
      #3955 TDC Pedestal run, aborted due to error
      #3956 TDC Pedestal run
    
    3.T1 has been repaird(broken wire) and put back together
    	-Please leave T1 chamber voltage off until 
    	further notice
    
      ZDC believed to be fixed
    
      Xterm for RHIC monitor replaced
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Wed Aug 01 2001 - 15:31:17 EDT